# BlueSuffix: Reinforced Blue Teaming for Vision-Language Models Against Jailbreak Attacks

### v1

@@ Abstract
Despite their superb multimodal capabilities, Vision-Language Models (VLMs) have been shown to be vulnerable to jailbreak attacks, which are inference-time attacks that induce the model to output harmful responses with tricky prompts. It is thus essential to defend VLMs against potential jailbreaks for their trustworthy deployment in real-world applications. In this work, we focus on black-box defense for VLMs against jailbreak attacks. Existing black-box defense methods are either unimodal or bimodal. Unimodal methods enhance either the vision or language module of the VLM, while bimodal methods robustify the model through text-image representation realignment. However, these methods suffer from two limitations: 1) they fail to fully exploit the cross-modal information, or 2) they degrade the model performance on benign inputs. To address these limitations, we propose a novel blue-team method BlueSuffix that defends the black-box target VLM against jailbreak attacks without compromising its performance. BlueSuffix includes three key components: 1) a visual purifier against jailbreak images, 2) a textual purifier against jailbreak texts, and 3) a blue-team suffix generator fine-tuned via reinforcement learning for enhancing cross-modal robustness. We empirically show on three VLMs (LLaVA, MiniGPT-4, and Gemini) and two safety benchmarks (MM-SafetyBench and RedTeam-2K) that BlueSuffix outperforms the baseline defenses by a significant margin. Our BlueSuffix opens up a promising direction for defending VLMs against jailbreak attacks. Correspondence to Xingjun Ma: <xingjunma@fudan.edu.cn>
@@ 1 Introduction
There has been a notable surge in research focusing on incorporating multimodal capabilities into Large Language Models (LLMs), leading to the emergence of Vision-Language Models (VLMs), such as OpenAI’s GPT-4o (Achiam et al., 2023) and Google’s Gemini 1.5 (Reid et al., 2024). VLMs leverage the combination of visual and textual modalities to perform a broad range of tasks, including image captioning and visual question answering, thereby extending the functionality of traditional LLMs. However, the integration of multi-modality introduces additional attack surfaces, bringing new security and safety challenges, particularly in their vulnerability to cross-modal jailbreak attacks that exploit maliciously crafted multimodal inputs to subvert the target VLM’s behaviors (Carlini et al., 2024; Bagdasaryan et al., 2023; Qi et al., 2023; Bailey et al., 2023; Gong et al., 2023; Wang et al., 2024; Fang et al., 2024; Ying et al., 2024). Addressing these vulnerabilities is thus critical for ensuring VLMs’ safe and reliable application in real-world scenarios.
Existing defense methods against VLM jailbreak attacks can be roughly divided into two types: 1) white-box defense that robustifies the VLM in the parameter space via adversarial training or fine-tuning, and 2) black-box defense that safeguards the input/output of the model in the prompt/response space using filters, detectors, or safety-driven system prompts. Arguably, black-box defense is more flexible and practical than white-box defense as it can protect the target VLM without accessing its parameters. In this paper, we focus on black-box defense against VLM jailbreak attacks.
@@
In this work, we focus on black-box defense against VLM jailbreaks and take a blue-team approach by training a blue-team suffix generator using reinforcement learning. Specifically, we propose a novel defense framework named BlueSuffix that leverages both unimodal and bimodal techniques to safeguard VLMs under a black-box defense setting, as illustrated in Figure 1. BlueSuffix has three key components: 1) a diffusion-based image purifier to defend the visual input against adversarial perturbations, 2) an LLM-based text purifier to rewrite the textual prompt following a certain template, enabling the target VLM to identify harmful content without altering its original meaning, and 3) an LLM-based blue-team suffix generator, which is fine-tuned from a lightweight language model using reinforcement learning to incorporate both visual and textual information for cross-modal robustness. In BlueSuffix, the image and text purifiers address the unimodal vulnerabilities of VLMs, while the blue-team suffix generator tackles the cross-modal vulnerabilities.
When training the blue-team suffix generator, we propose a novel cross-modal optimization strategy based on reinforcement learning to fine-tune a lightweight LLM. The optimization process takes the image and text purifiers into consideration to explore cross-model robustness. Specifically, it fine-tunes a GPT-2 based blue-team suffix generator to maximize the expected safety score given by an LLM-based judge (e.g., GPT-4o or Llama 3). The generated blue-team suffix does not affect the readability of the original textual prompt nor compromise the quality of the model’s response. During inference time, a defensive textual suffix will be generated (by the blue-team suffix generator), appended to the purified text input (by the text purifier), and fed into the target VLM along with the purified image input (by the image purifier). Unlike previous bimodal defenses, our approach does not focus on detecting malicious inputs. Instead, it mitigates malicious prompts through purification processes and the addition of blue-team suffixes. This makes it easier for the target model to recognize and respond to them correctly, making our method more practical for real-world scenarios.
@@
  • We propose a novel blue-team method, BlueSuffix, designed to protect black-box VLMs from generating harmful content in response to jailbreak prompts, particularly against universal adversarial visual triggers.
@@ 2 Related Work
VLMs are vision-integrated LLMs designed to process both visual and textual data, generating textual outputs for multimodal tasks. A typical VLM architecture comprises three key components: an image encoder, a text encoder, and a fusion module to integrate information from both encoders. For instance, MiniGPT-4 (Zhu et al., 2023) aligns visual data with a language model via a linear projection layer, connecting the pre-trained Vision Transformer (ViT) (Dosovitskiy, 2020) and Q-Former (Li et al., 2023a) to a frozen Vicuna model (Chiang et al., 2023). Similarly, LLaVA (Liu et al., 2024a) links the CLIP visual encoder (Radford et al., 2021) with the Vicuna model (Chiang et al., 2023) for general-purpose visual and language understanding. Building upon the pre-trained BLIP-2 models (Li et al., 2023a), InstructionBLIP (Dai et al., 2023) conducts a comprehensive study on vision-language instruction tuning and employs the Q-Former to synchronize visual features with the language model, thus boosting the model’s ability to interpret and respond to instruction-based queries.
@@
Jailbreak attacks aim to design malicious prompts that can bypass the safety mechanisms of an LLM or VLM to make it output harmful content. In the context of VLMs, jailbreak attacks are typically executed through carefully crafted malicious prompts that exploit vulnerabilities of the target model. Existing attack methods are either unimodal or bimodal. For unimodal attack, Zou et al. (2023) introduced a white-box method to optimize a universal adversarial suffix using the greedy coordinate gradient. Apart from universal adversarial suffixes, jailbreak can also be launched by template completion (Li et al., 2023b; Kang et al., 2024), prompt rewriting (Yuan et al., 2023; Yong et al., 2023), or LLM-based generation (Deng et al., 2024; Zeng et al., 2024a). The above methods were all initially designed for LLMs. Undoubtedly, jailbreak can also be achieved via adversarial images (Carlini et al., 2024; Niu et al., 2024). Subsequently, Qi et al. (2023) introduced a universal adversarial visual input. However, these methods are all unimodal attacks that fail to fully exploit the multimodal information in VLMs. Wang et al. (2024) employed dual optimization objectives to guide the generation of effective multimodal jailbreak prompts (i.e., chained texts and images). However, this attack only works in a white-box setting. Ying et al. (2024) proposed a Bi-Modal Adversarial Prompt Attack (BAP) to optimize query-agnostic universal adversarial perturbations (UAPs) and rewrite malicious textual prompts. BAP demonstrates universal attacking abilities across different scenarios.
@@
Accordingly, existing defenses against VLM jailbreak can also be categorized into unimodal and bimodal methods. For unimodal defense, white-box defense techniques can be applied to robustify the language model of VLM, for example instruction tuning (Bianchi et al., 2023; Deng et al., 2023), Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022; Siththaranjan et al., 2023), gradient analysis (Xie et al., 2024; Xu et al., 2024b), refinement (Kim et al., 2024; Zhang et al., 2024), and proxy defense (Zeng et al., 2024b; Struppek et al., 2024). While white-box defenses require full access to the model parameters, black-box defenses can protect the target model simply based on its inputs and outputs. Compared to white-box defenses, black-box defenses are often more flexible, lightweight, and effective. Existing black-box defenses for VLMs include prompt detection (Jain et al., 2023; Alon & Kamfonas, 2023; Liu et al., 2024c), prompt perturbation (Cao et al., 2023; Robey et al., 2023; Zhou et al., 2024; Liu et al., 2024b), and safety system prompt safeguards (Sharma et al., 2024; Zou et al., 2024; Zheng et al., 2024). The above-mentioned defense methods thus far are all language-based defenses. Apart from these methods, image denoising/purification methods can be applied to fix the jailbreak images. A well-known method is the DiffPure (Nie et al., 2022) which leverages a diffusion model to remove potential adversarial perturbations from the input images. However, this method only addresses the robustness of the visual modality.
Conversely, Jailguard (Zhang et al., 2023) trained a bimodal detector to identify malicious texts or images based on input mutation. Similar to Jailguard, CIDER (Xu et al., 2024a) leveraged cross-modal similarity between harmful queries and adversarial images to detect malicious inputs. While Jailguard’s performance heavily depends on the original alignment of VLMs, CIDER tends to negatively impact the model’s performance on benign queries. To address these limitations, in this work, we propose a novel blue-team framework BlueSuffix for black-box VLMs.
@@ 3.2 BlueSuffix
We utilize GPT-4o (Achiam et al., 2023) to achieve the above objective with a rewritten template. As GPT-4o is a commercial model, we also test the open-source model Llama-3-8B-Instruct (AI@Meta, 2024) as the text purifier. The results are referred to Appendix B which show that the prompts rewritten using LLaMA demonstrate a comparable performance to those by GPT-4o, in terms of both semantic expression and defense effectiveness.
@@
We denote the suffix generator as $\pi$, which receives a rewritten textual prompt $\hat{x}_{t}$ and generates a fixed-length suffix $x_{s}\sim\pi(\cdot|\hat{x}_{t})$. The suffix will be appended to the rewritten textual prompt as the final textual input as the target VLM. We denote the response of the target VLM as $y$ and leverage an LLM (GPT-4o or Llama-3-8B-Instruct) to judge the response. The output of the judge is a safety score which will then be used as the reward model $R(\cdot)$. The reward is either 1 or 0, i.e., $R(y)\in\{0,1\}$, with “1” representing benign response and “0” representing harmful response.
@@
$\max_{\pi}\mathbb{E}_{x_{s}\sim\pi(\cdot|\hat{x}_{t})}[R(y)-\beta D_{KL}(\pi(\cdot|\hat{x}_{t})\parallel\pi^{\text{ref}}(\cdot|\hat{x}_{t}))],$ (7)
@@
7:         Generate fixed-length suffix $x_{s}^{j}\sim\pi(\cdot|x_{t}^{j})$
@@
11:     Fine-tune the suffix generator $\pi=\arg\max_{\pi}\mathbb{E}_{x_{s}\sim\pi(\cdot|x_{t})}\left[R(y)-\beta D_{KL}\left(\pi(\cdot|x_{t})\parallel\pi^{\text{ref}}(\cdot|x_{t})\right)\right]$
@@ 4 Experiments
In this section, we evaluate our BlueSuffix defense on three VLMs and two safety benchmark datasets, focusing on its effectiveness, transferability, and robustness.
@@ 4.1 Experimental Setup
We test our defense on three VLMs, including two commonly used open-source large VLMs LLaVA (LLaVA-v1.5-7B) (Liu et al., 2024a) and MiniGPT-4 (Vicuna) (Zhu et al., 2023), and a commercial black-box VLM Gemini (gemini-1.5-flash) (Reid et al., 2024). We run our experiments on two popular safety benchmarks: MM-SafetyBench (Liu et al., 2023) and RedTeam-2K (Luo et al., 2024). MM-SafetyBench is a widely used safety benchmark dataset that consists of $1,680$ questions across 13 safety topics (unsafe scenarios) listed by OpenAI, such as privacy violation, fraudulent, and illegal activities. RedTeam-2K is a meticulously curated collection of $2,000$ harmful queries aimed at identifying alignment vulnerabilities in VLMs. It spans 16 safety policies and incorporates queries from 8 distinct sources. We evaluate the effectiveness of the defense methods on MM-SafetyBench and test its transferability to RedTeam-2K. We attack the target VLM using two types of attacks: 1) vanilla attack that directly inputs the jailbreak texts with the clean images into the model, and 2) a state-of-the-art bimodal attack BAP, which converts the clean images into jailbreak images via universal adversarial perturbation (UAP) while enhances the original jailbreak texts using ChatGPT.
@@
We compare our method with two black-box defense methods: DiffPure (Nie et al., 2022) and Safety Prompt Zheng et al. (2024). DiffPure is a general-purpose defense method that uses a diffusion model to remove the visual adversarial perturbations. DiffPure can be applied to purify any type of malicious input images, thus can be applied to purify jailbreak images. Safety Prompt inserts a defensive system prompt in front of the textual prompt, which acts as a “hint” to the target VLMs. We also combine DiffPure and Safety Prompt into a bimodal defense as a baseline. Note that we did not compare with Jailguard as it is a detection method that detects and rejects potential jailbreaks for the target model. Conversely, our defense weakens the malicious inputs such that the target model can identify the risk by itself and response robustly.
@@
We take the Attack Success Rate (ASR) as the primary performance metric. ASR quantifies the risk of the target model generating harmful content in the presence of jailbreak inputs. As the output of VLM are texts, we need a external judge to determine whether the response contains harmful content. Here, we use GPT-4o as the judge and design a system prompt to ask GPT-4o to classify the response: harmful vs. benign.
@@
For the image purifier, we directly adopt the denoising diffusion model released by DiffPure, as it has been shown to have high effectiveness and generality. For the text purifier, we test two LLM models: Llama-3-8B-Instruct (AI@Meta, 2024) and GPT-4o (Achiam et al., 2023)). The text purifier is instructed to rewrite the text input without altering its original meaning (the prompt template is provided in the Appendix C). The blue suffix generator is fine-tuned from a pre-trained GPT-2 using Proximal Policy Optimization (PPO) on hard jailbreak prompts crafted by the BAP attack (Ying et al., 2024) on the “financial advice” topic from the MM-SafetyBench. Please note that fine-tuned GPT-2 will be applied to other topics and other dataset (i.e., RedTeam-2K) to test its generalizability. The fine-tuning batch size is set to 32. The reward given by the LLM judge (i.e., GPT-4o) is “1” if the model’s response is benign, “0” otherwise. The fine-tuning can be stopped until the expected safety score exceeds 0.95, for about 300 epochs.

### v2

# Vision-Language Models Against Jailbreak Attacks

@@ Abstract
In this paper, we focus on black-box defense for VLMs against jailbreak attacks. Existing black-box defense methods are either unimodal or bimodal. Unimodal methods enhance either the vision or language module of the VLM, while bimodal methods robustify the model through text-image representation realignment. However, these methods suffer from two limitations: 1) they fail to fully exploit the cross-modal information, or 2) they degrade the model performance on benign inputs. To address these limitations, we propose a novel blue-team method BlueSuffix that defends target VLMs against jailbreak attacks without compromising its performance under black-box setting. BlueSuffix includes three key components: 1) a visual purifier against jailbreak images, 2) a textual purifier against jailbreak texts, and 3) a blue-team suffix generator using reinforcement fine-tuning for enhancing cross-modal robustness. We empirically show on four VLMs (LLaVA, MiniGPT-4, InstructionBLIP, and Gemini) and four safety benchmarks (Harmful Instruction, AdvBench, MM-SafetyBench, and RedTeam-2K) that BlueSuffix outperforms the baseline defenses by a significant margin. Our BlueSuffix opens up a promising direction for defending VLMs against jailbreak attacks. Code is available at https://github.com/Vinsonzyh/BlueSuffix.
@@ 1 Introduction
There has been a notable surge in research focusing on incorporating multimodal capabilities into Large Language Models (LLMs), leading to the emergence of Vision-Language Models (VLMs), such as OpenAI’s GPT-4o (Achiam et al., 2023), Google’s Gemini 1.5 (Reid et al., 2024) and DeepSeek’s DeepSeek-VL2 (Wu et al., 2024). VLMs leverage the combination of visual and textual modalities to perform a broad range of tasks, including image captioning and visual question answering, thereby extending the functionality of traditional LLMs. However, the integration of multi-modality introduces additional attack surfaces, bringing new security and safety challenges, particularly in their vulnerability to cross-modal jailbreak attacks that exploit maliciously crafted multimodal inputs to subvert the target VLM’s behaviors (Carlini et al., 2024; Bagdasaryan et al., 2023; Qi et al., 2024; Bailey et al., 2024; Gong et al., 2025; Wang et al., 2024; Fang et al., 2024; Ying et al., 2024). Addressing these vulnerabilities is thus critical for ensuring VLMs’ safe and reliable application in real-world scenarios.
Existing defense methods against VLM jailbreak attacks can be roughly divided into two types: 1) white-box defense that robustifies the VLM in the parameter space via adversarial training or fine-tuning, and 2) black-box defense that safeguards the input/output of the model in the prompt/response space using filters, detectors, or safety-driven system prompts. Arguably, black-box defense is more flexible and practical than white-box defense as it can protect the target VLM without accessing its parameters. In this paper, we focus on black-box defense.
@@
In this work, we focus on black-box defense and take a blue-team approach by training a suffix generator using reinforcement fine-tuning. Specifically, we propose a novel defense framework named BlueSuffix that leverages both unimodal and bimodal techniques to safeguard VLMs under a black-box defense setting, as illustrated in Figure 1. BlueSuffix has three key components: 1) a diffusion-based image purifier to defend the visual input against adversarial perturbations, 2) an LLM-based text purifier to rewrite the textual prompt following a certain template without altering its original meaning, and 3) an LLM-based blue-team suffix generator, which is fine-tuned from a lightweight language model using reinforcement learning to incorporate both visual and textual information for cross-modal robustness. In BlueSuffix, the image and text purifiers address the unimodal vulnerabilities, while the suffix generator tackles the cross-modal vulnerabilities.
When training the blue-team suffix generator, we propose a novel cross-modal optimization strategy based on reinforcement fine-tuning a lightweight LLM. The optimization process takes the image and text purifiers into consideration to explore cross-model robustness. Specifically, it fine-tunes a GPT-2 based blue-team suffix generator to maximize the expected safety score given by an LLM-based judge (e.g., GPT-4o or Llama 3). The generated blue-team suffix does not affect the readability of the original textual prompt nor compromise the quality of the model’s response. During inference time, a defensive textual suffix will be generated (by the blue-team suffix generator), appended to the purified text input (by the text purifier), and fed into the target VLM along with the purified image input (by the image purifier). Unlike previous bimodal defenses, our approach does not focus on detecting malicious inputs. Instead, it mitigates malicious prompts through purification processes and the addition of blue-team suffixes. This makes it easier for the target model to recognize and respond to them correctly, making our method more practical for real-world scenarios.
@@
  • We propose a novel blue-team framework, BlueSuffix, providing a plug-and-play, model-agnostic, and generic solution for blue-teaming VLMs under black-box setting, enabling seamless integration and extension of existing techniques.
@@ 2 Related Work
VLMs are vision-integrated LLMs designed to process both visual and textual data, generating textual outputs for multimodal tasks. A typical VLM architecture comprises three key components: an image encoder, a text encoder, and a fusion module to integrate information from both encoders. For instance, MiniGPT-4 (Zhu et al., 2024) aligns visual data with a language model via a linear projection layer, connecting the pre-trained Vision Transformer (ViT) (Dosovitskiy, 2020) and Q-Former (Li et al., 2023a) to a frozen Vicuna model (Chiang et al., 2023). Similarly, LLaVA (Liu et al., 2023) links the CLIP visual encoder (Radford et al., 2021) with the Vicuna model (Chiang et al., 2023) for general-purpose visual and language understanding. Building upon the pre-trained BLIP-2 models (Li et al., 2023a), InstructionBLIP (Dai et al., 2023) conducts a comprehensive study on vision-language instruction tuning and employs the Q-Former to synchronize visual features with the language model, thus boosting the model’s ability to interpret and respond to instruction-based queries.
@@
Jailbreak attacks aim to design malicious prompts that can bypass the safety mechanisms of an LLM or VLM to make it output harmful content. In the context of VLMs, jailbreak attacks are typically executed through carefully crafted malicious prompts that exploit vulnerabilities of the target model. Existing attack methods are either unimodal or bimodal. For unimodal attack, Zou et al. (2023) introduced a white-box method to optimize a universal adversarial suffix using the greedy coordinate gradient. Apart from universal adversarial suffixes, jailbreak can also be launched by template completion (Li et al., 2024; Kang et al., 2024), prompt rewriting (Yuan et al., 2024; Yong et al., 2023), or LLM-based generation (Deng et al., 2024; Zeng et al., 2024a). The above methods were all initially designed for LLMs. Undoubtedly, jailbreak can also be achieved via adversarial images (Carlini et al., 2024; Niu et al., 2024). Subsequently, Qi et al. (2024) introduced a universal adversarial visual input. However, these methods are all unimodal attacks that fail to fully exploit the multimodal information in VLMs. Wang et al. (2024) employed dual optimization objectives to guide the generation of effective multimodal jailbreak prompts (i.e., chained texts and images). However, this attack only works in a white-box setting. Ying et al. (2024) proposed a Bi-Modal Adversarial Prompt Attack (BAP) to optimize query-agnostic universal adversarial perturbations (UAPs) and rewrite malicious textual prompts. BAP demonstrates universal attacking abilities across different scenarios.
@@
Accordingly, existing defenses against VLM jailbreak can also be categorized into unimodal and bimodal methods. For unimodal defense, white-box defense techniques can be applied to robustify the language model of VLM, for example instruction tuning (Bianchi et al., 2024; Deng et al., 2023), Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022; Siththaranjan et al., 2024), gradient analysis (Xie et al., 2024; Xu et al., 2024b), refinement (Kim et al., 2024; Zhang et al., 2025), and proxy defense (Zeng et al., 2024b; Struppek et al., 2024). While white-box defenses require full access to the model parameters, black-box defenses can protect the target model simply based on its inputs and outputs. Compared to white-box defenses, black-box defenses are often more flexible, lightweight, and effective. Existing black-box defenses for VLMs include prompt detection (Jain et al., 2023; Alon & Kamfonas, 2023; Liu et al., 2024d), prompt perturbation (Cao et al., 2024; Robey et al., 2023; Zhou et al., 2024; Liu et al., 2024a), and safety system prompt safeguards (Sharma et al., 2024; Zou et al., 2024; Zheng et al., 2024). The above-mentioned defense methods thus far are all language-based defenses. Apart from these methods, image denoising/purification methods can be applied to fix the jailbreak images. A well-known method is the DiffPure (Nie et al., 2022) which leverages a diffusion model to remove potential adversarial perturbations from the input images. However, this method only addresses the robustness of the visual modality.
Conversely, Jailguard (Zhang et al., 2023) trained a bimodal detector to identify malicious texts or images based on input mutation. Similar to Jailguard, CIDER (Xu et al., 2024a) leveraged cross-modal similarity between harmful queries and adversarial images to detect malicious inputs. While Jailguard’s performance heavily depends on the original alignment of VLMs, CIDER tends to negatively impact the model’s performance on benign queries. To address these limitations, in this work, we propose a novel blue-team framework BlueSuffix for protecting VLMs under black-box setting.
@@ 3.2 BlueSuffix
We utilize GPT-4o (Achiam et al., 2023) to achieve the above objective with a rewritten template. As GPT-4o is a commercial model, we also test the open-source model Llama-3-8B-Instruct (AI@Meta, 2024) as the text purifier. The results are referred to Appendix A which show that the prompts rewritten using LLaMA demonstrate a comparable performance to those by GPT-4o, in terms of both semantic expression and defense effectiveness.
@@
We denote the suffix generator as $\pi$, which receives a rewritten textual prompt $\hat{x}_{t}$ and generates a fixed-length suffix $x_{\text{suffix}}\sim\pi(\cdot|\hat{x}_{t})$. The suffix will be appended to the rewritten textual prompt as the final textual input as the target VLM. We denote the response of the target VLM as $y$ and leverage an LLM (GPT-4o or Llama-3-8B-Instruct) to judge the response. The output of the judge is a safety score which will then be used as the reward model $R(\cdot)$. The reward is either 1 or 0, i.e., $R(y)\in\{0,1\}$, with “1” representing benign response and “0” representing harmful response.
@@
$\max_{\pi}\mathbb{E}_{x_{\text{suffix}}\sim\pi(\cdot|\hat{x}_{t})}[R(y)-\beta D_{KL}(\pi(\cdot|\hat{x}_{t})\parallel\pi^{\text{ref}}(\cdot|\hat{x}_{t}))],$ (7)
@@
7:         Generate fixed-length suffix $x_{\text{suffix}}^{j}\sim\pi(\cdot|x_{t}^{j})$
@@
11:     Fine-tune the suffix generator $\pi=\arg\max_{\pi}\mathbb{E}_{x_{\text{suffix}}\sim\pi(\cdot|x_{t})}\left[R(y)-\beta D_{KL}\left(\pi(\cdot|x_{t})\parallel\pi^{\text{ref}}(\cdot|x_{t})\right)\right]$
@@ 4 Experiments
In this section, we evaluate our BlueSuffix defense on four VLMs and four safety benchmark datasets, focusing on its effectiveness, transferability, and robustness.
@@ 4.1 Experimental Setup
We test our defense on four VLMs, including three commonly used open-source large VLMs LLaVA (LLaVA-v1.5-7B) (Liu et al., 2023), MiniGPT-4 (Vicuna) (Zhu et al., 2024), and InstructionBLIP (Vicuna) (Dai et al., 2023) as well as a commercial black-box VLM Gemini (gemini-1.5-flash) (Reid et al., 2024). We run our experiments on four popular safety benchmarks: AdvBench (Zou et al., 2023), MM-SafetyBench (Liu et al., 2024c), RedTeam-2K (Luo et al., 2024) and Harmful_Instructions (Qi et al., 2024). Detailed introductions of the safety benchmarks are provided in the Appendix E.
Jailbreak Methods
We attack the target VLM using six types of attacks, including two image-based attacks: Visual Adversarial Attacks (VAA) (Qi et al., 2024) and image Jailbreaking Prompt (imgJP) (Niu et al., 2024), two text-based attacks: Greedy Coordinate Gradient (GCG) (Zou et al., 2023) and AutoDAN (Liu et al., 2024b) as well as two bimodal attacks: Vanilla Attack and BAP Attack (Ying et al., 2024). Detailed introductions to the jailbreak methods and the corresponding datasets used are provided in the Appendix F.
@@
We compare our method with six defense methods, including 1) DiffPure (Nie et al., 2022), 2) Safety Prompt (Zheng et al., 2024), and 3) the combination of DiffPure and Safety Prompt 4) Robust Refusal Dynamic Defense (R2D2) (Mazeika et al., 2024), 5) Continuous Adversarial Training (CAT) (Xhonneux et al., 2024), and 6) VLGuard (Zong et al., 2024). The technical details of baseline defenses are in Appendix G.
@@
We take the Attack Success Rate (ASR) as the primary performance metric. ASR quantifies the risk of the target model generating harmful content in the presence of jailbreak inputs. As the output of VLM are texts, we need a external judge to determine whether the response contains harmful content. Here, we use GPT-4o as the judge and design a system prompt to ask GPT-4o to classify the response: harmful vs. benign. We also employed the Perspective API to score the toxicity of LLaVA model responses under vanilla attacks, BAP attacks, and when defended by BlueSuffix against BAP attacks on the MM-SafetyBench dataset. The Perspective API provides scores ranging from 0 to 1, with higher scores indicating a greater likelihood that the comment is considered toxic. The experimental results are shown in Appendix H.
@@
For the image purifier, we directly adopt the denoising diffusion model released by DiffPure, as it has been shown to have high effectiveness and generality. For the text purifier, we test two LLM models: Llama-3-8B-Instruct (AI@Meta, 2024) and GPT-4o (Achiam et al., 2023)). The text purifier is instructed to rewrite the text input without altering its original meaning (the prompt template is provided in the Appendix C). The blue suffix generator is fine-tuned from a pre-trained GPT-2 using Proximal Policy Optimization (PPO) (Schulman et al., 2017) on hard jailbreak prompts crafted by the BAP attack (Ying et al., 2024) on all 13 jailbreak topics from the MM-SafetyBench. Please note that fine-tuned GPT-2 will be applied to defend other attacks (ImgJP, VAA, GCG, and AutoDAN) and other datasets (RedTeam-2K, AdvBench, and Harmful_Instructions) to test its generalizability. The fine-tuning batch size is set to 32. The reward given by the LLM judge (i.e., GPT-4o, the judge template is provided in the Appendix D) is “1” if the model’s response is benign, “0” otherwise. The fine-tuning can be stopped until the expected safety score exceeds 0.95, for about 300 epochs.
Table 1: The ASR (%) achieved by different defense methods against various attacks (first column). A lower ASR denotes better defense performance. The format “Attack (Dataset)” denotes evaluating Attack using Dataset. For target VLMs, the format “Model A (Model B)” means defending black-box Model A against jailbreak images generated by BAP on white-box Model B (as UAP generation is white-box). It is important to note that VLGuard is only available for the LLaVA model and the R2D2 and CAT are white-box defense method, so they are only tested on open-source models..
Attack (Dataset) Model No Defense DiffPure System Prompt DiffPure + System Prompt R2D2 CAT VLGuard BlueSuffix VAA (Harmful_Instructions) LLaVA-v1.5-7B 57.50 42.50 50.00 35.00 17.50 5.00 10.00 0.00 MiniGPT-4 47.50 40.00 27.50 20.00 5.00 10.00 - 0.00 InstructionBLIP 42.50 37.50 37.50 17.50 15.00 5.00 - 0.00 Gemini-1.5-flash(MiniGPT-4) 10.00 0.00 10.00 12.50 - - - 0.00 Gemini-1.5-flash(LLaVA) 2.50 0.00 2.50 5.00 - - - 0.00 Gemini-1.5-flash(InstructionBLIP) 5.00 0.00 2.50 5.00 - - - 0.00 ImgJP (AdvBench) LLaVA-v1.5-7B 75.00 45.00 34.00 19.00 63.00 17.00 57.00 0.00 MiniGPT-4 66.00 41.00 24.00 16.00 17.00 21.00 - 0.00 InstructionBLIP 45.00 32.00 27.00 17.00 23.00 13.00 - 0.00 Gemini-1.5-flash 8.00 4.00 5.00 2.00 - - - 0.00 GCG (AdvBench) LLaVA-v1.5-7B 60.00 59.00 21.00 21.00 35.00 7.00 58.00 0.00 MiniGPT-4 46.00 44.00 16.00 15.00 8.00 15.00 - 0.00 InstructionBLIP 58.00 58.00 22.00 21.00 30.00 15.00 - 0.00 Gemini-1.5-flash 7.00 7.00 2.00 1.00 - - - 0.00 AutoDAN (AdvBench) LLaVA-v1.5-7B 80.00 80.00 27.00 25.00 26.00 30.00 71.00 0.00 MiniGPT-4 59.00 58.00 19.00 19.00 28.00 11.00 - 0.00 InstructionBLIP 60.00 61.00 22.00 20.00 45.00 26.00 - 0.00 Gemini-1.5-flash 8.00 7.00 2.00 1.00 - - - 0.00 BAP Attack (MM-SafetyBench) LLaVA-v1.5-7B 61.02 32.47 28.36 11.84 46.55 41.85 21.67 4.65 MiniGPT-4 62.26 22.83 21.42 16.07 34.94 40.89 - 9.37 InstructionBLIP 58.48 25.25 22.68 19.29 35.00 25.89 - 6.78 Gemini-1.5-flash(LLaVA) 40.98 2.73 2.22 2.92 - - - 0.47 Gemini-1.5-flash(MiniGPT-4) 41.07 1.43 1.77 2.10 - - - 0.30 Gemini-1.5-flash(InstructionBLIP) 40.71 2.62 2.08 2.80 - - - 0.42